OpenStudy (anonymous):

2y-1/4y = 4/6 , solve for y?

OpenStudy (mertsj):

Cross multiply

OpenStudy (mertsj):

\[6(2y-1)=4(4y)\]

OpenStudy (mertsj):

\[12y-6=16y\]

OpenStudy (mertsj):

\[-6=4y\]

OpenStudy (mertsj):

\[-\frac{3}{2}=y\]

OpenStudy (anonymous):

ahh okok I see I see. and sorry I think I posted this twice.